HETAV PATEL

Grade 11,

What is the potential at the apex of the cone (hollow and solid)?

What is the potential at the apex of the cone (hollow and solid)?

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,Wor is defined in electrostatic as:
Work done  = potential energy in bringing the charge from infinity to apex
dU = 1/4πɛ .dQq/x
U =∫du   (Integeral of u)
   =q/4πɛ x2Q/L∫xdx/x
W =Qq/2πɛL.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)